Problem 7-18 Comparing Investment Criteria

Consider the following cash flows of two mutually exclusive projects for A–Z Motorcars. Assume the discount rate for both projects is 11 percent.

Year AZM
Mini-SUV
AZF
Full-SUV
0 –$ 545,000 –$ 895,000
1 339,000 369,000
2 218,000 458,000
3 169,000 309,000


a. What is the payback period for each project? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Payback period
AZM Mini-SUV years
AZF Full-SUV years


b. What is the NPV for each project? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

NPV
AZM Mini-SUV $
AZF Full-SUV $


c. What is the IRR for each project? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)

IRR
AZM Mini-SUV %
AZF Full-SUV %
Answer #1

Answer:

Pay Back Period:

AZM MINI - SUV = 1.06 Years

AZM FULL - SUV = 2.78 Years

According to payback period AZM MINI -SUV Should be accepted . (Lower Pay Back Period)

NPV:

AZM MINI - SUV = $ 60910.44

AZM FULL - SUV = $ 35093.64

According to NPV Method AZM MINI -SUV Should be accepted . (Higher NPV)

IRR :

AZM MINI - SUV = 18.22%

AZM FULL - SUV = 13.30%

According to IRR Method AZM MINI -SUV Should be accepted . (Higher IRR)

Calculations are given below:

(a) Pay Back Period:

Year AZM MINI - SUV Cumulative AZF FULL - SUV Cumulative
0 -545000 -895000
1 339000 339000 369000 369000
2 218000 557000 458000 827000
3 169000 726000 309000 1136000

Pay Back (AZM) = 1+(557000-545000)/218000 = 1.06

Pay Back (AZF) = 2+(1136000-895000)/309000 = 2.78

(b) NPV :

NPV - AZM MINI - SUV

Year AZM MINI - SUV PV Factor PV
0 -545000 1.00 -545000
1 339000 0.90 305405.4054
2 218000 0.81 176933.6904
3 169000 0.73 123571.3434
NPV 60910.44

NPV - AZF FULL - SUV

Year AZM MINI - SUV PV Factor PV
0 -895000 1.00 -895000
1 369000 0.90 332432.4324
2 458000 0.81 371723.0744
3 309000 0.73 225938.1368
NPV 35093.64

(c) IRR:

IRR - AZM MINI - SUV

Year AZM MINI - SUV PV Factor @ 18% PV PV Factor @ 19% PV
0 -545000 1.00 -545000 1.00 -545000
1 339000 0.85 287288.1356 0.84 284873.9
2 218000 0.72 156564.2057 0.71 153943.9
3 169000 0.61 102858.6175 0.59 100287.3
NPV (1) 1710.96 NPV (2) -5894.85

IRR = LDR + (NPV1)* (HDR-LDR)/(NPV1 - NPV2)

LDR = 18%

HDR = 19%

IRR = 18 + (19-18)* (1710.96)/(1710.96+5894.85) = 18.22%

IRR - AZF FULL - SUV

Year AZM MINI - SUV PV Factor @ 13% PV PV Factor @ 14% PV
0 -895000 1.00 -895000 1.00 -895000
1 369000 0.88 326548.6726 0.88 323684.2
2 458000 0.78 358681.181 0.77 352416.1
3 309000 0.69 214152.5001 0.67 208566.2
NPV (1) 4382.35 NPV (2) -10333.5

LDR = 13%

HDR = 14%

IRR = 13 + (4382.35)* (14-13)/ (4382.35+10333.5) = 13.30%